# CodeNER: Code Prompting for Named Entity Recognition

### v3

@@ Abstract
Recent studies have explored various approaches for treating candidate named entity spans as both source and target sequences in named entity recognition (NER) by leveraging large language models (LLMs). Although previous approaches have successfully generated candidate named entity spans with suitable labels, they rely solely on input context information when using LLMs, particularly, ChatGPT. However, NER inherently requires capturing detailed labeling requirements with input context information. To address this issue, we propose a novel method that leverages code-based prompting to improve the capabilities of LLMs in understanding and performing NER. By embedding code within prompts, we provide detailed BIO schema instructions for labeling, thereby exploiting the ability of LLMs to comprehend long-range scopes in programming languages. Experimental results demonstrate that the proposed code-based prompting method outperforms conventional text-based prompting on ten benchmarks across English, Arabic, Finnish, Danish, and German datasets, indicating the effectiveness of explicitly structuring NER instructions. We also verify that combining the proposed code-based prompting method with the chain-of-thought prompting further improves performance. Corresponding author.Our code and dataset are available at https://github.com/HanSungwoo/CodeNER.
@@
Sungwoo Han1, Jingun Kwon1, Hidetaka Kamigaito2, and Manabu Okumura3 1Chungnam National University, 2Nara Institute of Science and Technology (NAIST) 3Institute of Science Tokyo 77sungwhan@o.cnu.ac.kr jingun.kwon@cnu.ac.kr kamigaito.h@is.naist.jp oku@pi.titech.ac.jp
@@ 1 Introduction
Named entity recognition (NER) is a task of detecting and classifying named entities in a given text, such as person, location, and miscellaneous entities. Traditionally, NER has been treated as a sequential labeling problem, e.g., with the BIO output schema, where a learned model classifies each token in a sequence from a predefined label set carreras-etal-2003-learning; lample-etal-2016-neural; ma-hovy-2016-end; wan-etal-2022-nested. These schemata indicate that each token in the input should be aligned with the BIO tag by capturing entity boundaries. Incorporating conditional random field (CRF) networks, which explicitly capture broader context information of input, has led to leading performance 10.5555/645530.655813; finkel-etal-2005-incorporating; liu-etal-2011-recognizing; lample-etal-2016-neural; jie-lu-2019-dependency; kruengkrai-etal-2020-improving; xu-etal-2021-better.
Recently, large language models (LLMs) have achieved great success in various natural language processing (NLP) tasks by leveraging in-context learning with prompts and exploiting their zero-shot task-solving abilities without the need to update model parameters neuips2022; weifinetuned; 51115. There are two main types of LLMs: closed and open. While closed models, such as ChatGPT, which is based on GPT-4, [^1] https://chat.openai.com/ have shown strong reasoning abilities, open models, such as Llama-3 and Phi-3, offer flexibility to access and modify their architectures llama3modelcard; abdin2024phi3technicalreporthighly.
Motivated by these recent advancements, the zero-shot reasoning abilities of LLMs, specifically GPT, in NER have been explored xie-etal-2023-empirical; 10.1109/JCDL57899.2023.00034; xie-etal-2024-self; 10.1007/978-3-031-72437-4_22. Despite the success of using LLMs, however, previous work has relied on two-stage approaches, such as majority voting or constructing self-annotated datasets, for few-shot techniques with text-based prompting. Text-based prompting faces challenges in NER, primarily due to the inherent gap between the text-in-text-out schema of LLMs and the text-in-span-out nature of the task 10.5555/AAI28114605. This mismatch makes it difficult to handle the sequential aspect of the BIO output schema, which is integral to NER, in LLMs and can result in poor performance in identifying entity boundaries ding-etal-2024-rethinking. Although various prompting methods, such as chain-of-thought (CoT) han2023informationextractionsolvedchatgpt; xie-etal-2023-empirical, have been introduced to improve reasoning cot, challenges still remain in bridging this gap.
@@ 2 Related Work
Named Entity Recognition. NER is the task of identifying and annotating named entities in a given text from a predefined label set. Traditionally, NER has been approached as a sequential labeling task carreras-etal-2003-learning; lample-etal-2016-neural. Early efforts employed LSTM networks to represent token embeddings and subsequently integrated conditional random field (CRF) layers with CNNs to consider broader context information collobert. In addition, hybrid models that combine LSTM and CNNs have been investigated to leverage character-level information chiu-nichols-2016-named; ma-hovy-2016-end. Another approach formulates the NER task as span prediction, where learned models simultaneously detect span boundaries and annotate named entities stratos-2017-entity; li-etal-2020-unified; ouchi-etal-2020-instance; fu-etal-2021-spanner. Recent work has extended the span-based approaches to better capture nested entities, which can appear when one named entity is contained within another shen-etal-2021-locate; tan2021sequencetoset; ding-etal-2022-ask. Instead of formulating NER as a sequential labeling task, generative methods have been introduced, that utilize a Seq2Seq framework by considering candidate named entity spans as both source and target sequences cui-etal-2021-template; Fei_Ji_Li_Liu_Ren_Li_2021; yan-etal-2021-unified-generative. While supervised methods have achieved saturated performance on the standard NER benchmarks, they require extensive labeled datasets chen-etal-2021-data; zhou-etal-2022-melm and domain adaptation methods zhang-etal-2021-pdaln; yang-etal-2022-factmix; xu-etal-2023-improving.
@@
In-context Learning. LLMs have achieved significant success in various NLP tasks in zero-shot settings by leveraging specifically formatted instruction-based inputs Radford2019LanguageMA; NEURIPS2020_1457c0d6; chowdhery2022palm. Different from the fine-tuning setting, which requires additional parameter updates for learning to solve tasks, the zero-shot setting leverages in-context learning to perform tasks without any additional adjustments of the parameters of a pre-trained model. Owing to its flexibility, numerous efforts have been made to investigate better instructions for enhancing reasoning abilities, including few-shot, directional stimulus, and CoT learning liu-etal-2022-generated; zhu2023multilingual; cot; li2023guiding. Recently, the use of LLMs in NER has gained attention due to their remarkable reasoning abilities. wang2023gptnernamedentityrecognition considered a retrieval-augmented generation approach by leveraging entity-level information using a fine-tuned NER model to search exemplars. xie-etal-2023-empirical explored the reasoning abilities of ChatGPT by considering few-shot techniques based on a two-stage majority voting. xie-etal-2024-self demonstrated that self-annotating an unlabeled dataset using ChatGPT further improves performance with the two-stage majority voting. Additionally, instruction-based fine-tuning with LLMs and evaluating unseen labels on new datasets using different prompts have also been studied sainz2024gollie; ding-etal-2024-rethinking; zamai2024lessinstructmoreenriching.
@@ 3 CodeNER
Unlike text-based prompting methods, which rely solely on natural language instructions for NER, our CodeNER is a code-based prompting method. Figures 1 (a) and (b) show text-based prompts and code-based prompts in Python, respectively. CodeNER explicitly encourages models to consider the BIO output schema and broader context information of input via prompts written in a programming language-style. This approach is critical for improving performance in NER tasks since it enables the model to better capture entity boundaries ding-etal-2024-rethinking than existing methods using LLMs that consider only context information liu-etal-2011-recognizing; kruengkrai-etal-2020-improving; xu-etal-2021-better. Python code-based prompts provide a clear and human-readable structure while overcoming several challenges, such as potential misinterpretation and variability in natural languages wang-etal-2023-code4struct; LiSTYWHQ23; sainz2024gollieannotationguidelinesimprove.
@@ 4.1 Experimental Settings
Datasets. We used ten benchmark datasets. Table 1 provides the statistics of the NER test datasets. For the English datasets, we used CoNLL03 tjong-kim-sang-de-meulder-2003-introduction, FIN salinas-alvarado-etal-2015-domain, WNUT-17 derczynski-etal-2017-results, the MIT movie review semantic corpus, and the MIT restaurant review. [^3] The MIT NER datasets are publicly available at https://groups.csail.mit.edu/sls/downloads/. For the Arabic datasets, we used ArabicDatasets A and B bari19. For the Finnish dataset, we used FinnishDataset A Ruokolainen_2019. The German and Danish datasets are SwissNER vamvas-etal-2023-swissbert and DaNE hvingelby-etal-2020-dane, respectively.
We transformed the traditional NER datasets into a programming language-style format. Note that, when considering languages other than English, we used Google Translate to translate the “ner_tags_labels” descriptions in both CodeNER and Vanilla prompts. [^4] https://translate.google.com/ Different from previous work that sampled only a small portion of datasets xie-etal-2023-empirical; xie-etal-2024-self, we used the entire datasets to ensure a comprehensive evaluation.
Implementation Details. For closed models, we employed the recent ChatGPT (GPT-4 and GPT-4 Turbo) models with their official APIs. For open models, we employed open-source Llama-3-8B and Phi-3-mini-128k-instruct as the backbone models for our experiments. Following previous work, we used F1 scores across all datasets based on the model-generated outputs cui-etal-2021-template; xie-etal-2023-empirical by considering the traditional evaluation method of BIO tagging hammerton-2003-named; ma-hovy-2016-end; chen-etal-2020-local. All experiments for the open models were conducted using an NVIDIA RTX A6000 GPU. For the decoding step, we set the $\rm do_{sample}$ option to False. [^5] All hyperparameter settings are described in Appendix A.
Baseline. We compared the proposed CodeNER to vanilla text-based prompting NER, shown in Figure 1 (a) xie-etal-2023-empirical in zero-shot settings. [^6] We converted model-generated outputs to the BIO format for evaluation.
@@ 5.1 Open Models
We evaluated the proposed CodeNER with the open models, Phi-3 and Llama-3. We additionally compared our CodeNER with the prompts from zero-shot NER models, GoLLIE and GNER, which are partial code-based prompts that focus on context information and BIO output schema without code, respectively sainz2024gollie; ding-etal-2024-rethinking. Table 4 shows the $F_{1}$ scores, including the detailed performance for each label. For both models, the proposed CodeNER consistently outperformed Vanilla with an average improvement of 19.54, 5.81, and 2.45 points in the $F_{1}$ scores.
@@ 5.4 Combination with Chain-of-thought
We evaluated the proposed CodeNER by incorporating the CoT reasoning instruction, “Let’s think step by step”, at the beginning of the CodeNER instruction. Although previous text-based vanilla prompting methods have been also explored to incorporate CoT in NER han2023informationextractionsolvedchatgpt; xie-etal-2023-empirical, the proposed approach, CodeNER, presents a different perspective. It encourages LLMs to enhance their understanding of long-range scopes using programming language-style prompts. Table 6 shows the results. The integration of CoT with CodeNER demonstrated the potential to improve performance in zero-shot settings by leveraging structured programming language-style prompts.
@@ 5.5 Different Programming Languages
Interestingly, the $F_{1}$ scores with CodeNER remain consistent between the Python and C++ instructions on CoNLL03. We think this suggests that dataset characteristics may have a more significant impact than the choice of the programming language. The comparable results between Python and C++ suggest that the choice of the programming language has a minimal impact on the effectiveness of CodeNER when used on datasets where the model’s existing capabilities are strong. We additionally considered using the Java programming language for CodeNER. However, the performance was nearly zero in terms of the F1 scores, which indicates LLMs lack inherent knowledge of Java in their pre-training haroon2025accuratelylargelanguagemodels.

### v4

@@ Abstract
Recent studies have explored various approaches for treating candidate named entity spans as both source and target sequences in named entity recognition (NER) by leveraging large language models (LLMs). Although previous approaches have successfully generated candidate named entity spans with suitable labels, they rely solely on input context information when using LLMs, particularly, ChatGPT. However, NER inherently requires capturing detailed labeling requirements with input context information. To address this issue, we propose a novel method that leverages code-based prompting to improve the capabilities of LLMs in understanding and performing NER. By embedding code within prompts, we provide detailed BIO schema instructions for labeling, thereby exploiting the ability of LLMs to comprehend long-range scopes in programming languages. Experimental results demonstrate that the proposed code-based prompting method outperforms conventional text-based prompting on ten benchmarks across English, Arabic, Finnish, Danish, and German datasets, indicating the effectiveness of explicitly structuring NER instructions. We also verify that combining the proposed code-based prompting method with the chain-of-thought prompting further improves performance. [^1] Our code and dataset are available at https://github.com/XXXX.
@@
Sungwoo Han1†, Hyeyeon Kim1†, Jingun Kwon1, Hidetaka Kamigaito2, and Manabu Okumura3 1Chungnam National University, 2Nara Institute of Science and Technology (NAIST) 3Institute of Science Tokyo {77sungwhan,hyk22}@o.cnu.ac.kr jingun.kwon@cnu.ac.kr kamigaito.h@is.naist.jp oku@pi.titech.ac.jp
@@ 1 Introduction
Named entity recognition (NER) is a task of detecting and classifying named entities in a given text, such as person, location, and miscellaneous entities. Traditionally, NER has been treated as a sequential labeling problem, e.g., with the BIO output schema, where a learned model classifies each token in a sequence from a predefined label set Carreras et al. (2003); Lample et al. (2016); Ma and Hovy (2016); Wan et al. (2022). These schemata indicate that each token in the input should be aligned with the BIO tag by capturing entity boundaries. Incorporating conditional random field (CRF) networks, which explicitly capture broader context information of input, has led to leading performance Lafferty et al. (2001); Finkel et al. (2005); Liu et al. (2011); Lample et al. (2016); Jie and Lu (2019); Kruengkrai et al. (2020); Xu et al. (2021).
Recently, large language models (LLMs) have achieved great success in various natural language processing (NLP) tasks by leveraging in-context learning with prompts and exploiting their zero-shot task-solving abilities without the need to update model parameters Ouyang et al. (2022); Wei et al. (2022); Cohen et al. (2022). There are two main types of LLMs: closed and open. While closed models, such as ChatGPT, which is based on GPT-4, [^2] https://chat.openai.com/ have shown strong reasoning abilities, open models, such as Llama-3 and Phi-3, offer flexibility to access and modify their architectures AI@Meta (2024); Abdin et al. (2024).
Motivated by these recent advancements, the zero-shot reasoning abilities of LLMs, specifically GPT, in NER have been explored Xie et al. (2023); González-Gallardo et al. (2024a); Xie et al. (2024); González-Gallardo et al. (2024b). Despite the success of using LLMs, however, previous work has relied on two-stage approaches, such as majority voting or constructing self-annotated datasets, for few-shot techniques with text-based prompting. Text-based prompting faces challenges in NER, primarily due to the inherent gap between the text-in-text-out schema of LLMs and the text-in-span-out nature of the task Chen (2018). This mismatch makes it difficult to handle the sequential aspect of the BIO output schema, which is integral to NER, in LLMs and can result in poor performance in identifying entity boundaries Ding et al. (2024). Although various prompting methods, such as chain-of-thought (CoT) Han et al. (2023); Xie et al. (2023), have been introduced to improve reasoning Wei et al. (2024), challenges still remain in bridging this gap.
@@ 2 Related Work
Named Entity Recognition. NER is the task of identifying and annotating named entities in a given text from a predefined label set. Traditionally, NER has been approached as a sequential labeling task Carreras et al. (2003); Lample et al. (2016). Early efforts employed LSTM networks to represent token embeddings and subsequently integrated conditional random field (CRF) layers with CNNs to consider broader context information Collobert et al. (2011). In addition, hybrid models that combine LSTM and CNNs have been investigated to leverage character-level information Chiu and Nichols (2016); Ma and Hovy (2016). Another approach formulates the NER task as span prediction, where learned models simultaneously detect span boundaries and annotate named entities Stratos (2017); Li et al. (2020); Ouchi et al. (2020); Fu et al. (2021). Recent work has extended the span-based approaches to better capture nested entities, which can appear when one named entity is contained within another Shen et al. (2021); Tan et al. (2021); Ding et al. (2022). Instead of formulating NER as a sequential labeling task, generative methods have been introduced, that utilize a Seq2Seq framework by considering candidate named entity spans as both source and target sequences Cui et al. (2021); Fei et al. (2021); Yan et al. (2021). While supervised methods have achieved saturated performance on the standard NER benchmarks, they require extensive labeled datasets Chen et al. (2021); Zhou et al. (2022) and domain adaptation methods Zhang et al. (2021); Yang et al. (2022); Xu et al. (2023).
@@
In-context Learning. LLMs have achieved significant success in various NLP tasks in zero-shot settings by leveraging specifically formatted instruction-based inputs Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2022). Different from the fine-tuning setting, which requires additional parameter updates for learning to solve tasks, the zero-shot setting leverages in-context learning to perform tasks without any additional adjustments of the parameters of a pre-trained model. Owing to its flexibility, numerous efforts have been made to investigate better instructions for enhancing reasoning abilities, including few-shot, directional stimulus, and CoT learning Liu et al. (2022); Zhu et al. (2023); Wei et al. (2024); Li et al. (2023b). Recently, the use of LLMs in NER has gained attention due to their remarkable reasoning abilities. Wang et al. (2023a) considered a retrieval-augmented generation approach by leveraging entity-level information using a fine-tuned NER model to search exemplars. Xie et al. (2023) explored the reasoning abilities of ChatGPT by considering few-shot techniques based on a two-stage majority voting. Xie et al. (2024) demonstrated that self-annotating an unlabeled dataset using ChatGPT further improves performance with the two-stage majority voting. Additionally, instruction-based fine-tuning with LLMs and evaluating unseen labels on new datasets using different prompts have also been studied Sainz et al. (2024a); Ding et al. (2024); Zamai et al. (2024).
@@ 3 CodeNER
Unlike text-based prompting methods, which rely solely on natural language instructions for NER, our CodeNER is a code-based prompting method. Figures 1 (a) and (b) show text-based prompts and code-based prompts in Python, respectively. CodeNER explicitly encourages models to consider the BIO output schema and broader context information of input via prompts written in a programming language-style. This approach is critical for improving performance in NER tasks since it enables the model to better capture entity boundaries Ding et al. (2024) than existing methods using LLMs that consider only context information Liu et al. (2011); Kruengkrai et al. (2020); Xu et al. (2021). Python code-based prompts provide a clear and human-readable structure while overcoming several challenges, such as potential misinterpretation and variability in natural languages Wang et al. (2023b); Li et al. (2023a); Sainz et al. (2024b); Zeng et al. (2025).
@@ 4.1 Experimental Settings
Datasets. We used ten benchmark datasets. Table 1 provides the statistics of the NER test datasets. For the English datasets, we used CoNLL03 Tjong Kim Sang and De Meulder (2003), FIN Salinas Alvarado et al. (2015), WNUT-17 Derczynski et al. (2017), the MIT movie review semantic corpus, and the MIT restaurant review. [^4] The MIT NER datasets are publicly available at https://groups.csail.mit.edu/sls/downloads/. For the Arabic datasets, we used ArabicDatasets A and B Bari et al. (2020). For the Finnish dataset, we used FinnishDataset A Ruokolainen et al. (2019). The German and Danish datasets are SwissNER Vamvas et al. (2023) and DaNE Hvingelby et al. (2020), respectively.
We transformed the traditional NER datasets into a programming language-style format. Note that, when considering languages other than English, we used Google Translate to translate the “ner_tags_labels” descriptions in both CodeNER and Vanilla prompts. [^5] https://translate.google.com/ Different from previous work that sampled only a small portion of datasets Xie et al. (2023, 2024), we used the entire datasets to ensure a comprehensive evaluation.
Implementation Details. For closed models, we employed the recent ChatGPT (GPT-4 and GPT-4 Turbo) models with their official APIs. For open models, we employed open-source Llama-3-8B and Phi-3-mini-128k-instruct as the backbone models for our experiments. Following previous work, we used F1 scores across all datasets based on the model-generated outputs Cui et al. (2021); Xie et al. (2023) by considering the traditional evaluation method of BIO tagging Hammerton (2003); Ma and Hovy (2016); Chen et al. (2020). All experiments for the open models were conducted using an NVIDIA RTX A6000 GPU. For the decoding step, we set the $\rm do_{sample}$ option to False. [^6] All hyperparameter settings are described in Appendix A.
Baseline. We compared the proposed CodeNER to vanilla text-based prompting NER, shown in Figure 1 (a) Xie et al. (2023) in zero-shot settings. [^7] We converted model-generated outputs to the BIO format for evaluation.
@@ 5.1 Open Models
We evaluated the proposed CodeNER with the open models, Phi-3 and Llama-3. We additionally compared our CodeNER with the prompts from zero-shot NER models, GoLLIE and GNER, which are partial code-based prompts that focus on context information and BIO output schema without code, respectively Sainz et al. (2024a); Ding et al. (2024). Table 4 shows the $F_{1}$ scores, including the detailed performance for each label. For both models, the proposed CodeNER consistently outperformed Vanilla with an average improvement of 19.54, 5.81, and 2.45 points in the $F_{1}$ scores.
@@ 5.4 Combination with Chain-of-thought
We evaluated the proposed CodeNER by incorporating the CoT reasoning instruction, “Let’s think step by step”, at the beginning of the CodeNER instruction. Although previous text-based vanilla prompting methods have been also explored to incorporate CoT in NER Han et al. (2023); Xie et al. (2023), the proposed approach, CodeNER, presents a different perspective. It encourages LLMs to enhance their understanding of long-range scopes using programming language-style prompts. Table 6 shows the results. The integration of CoT with CodeNER demonstrated the potential to improve performance in zero-shot settings by leveraging structured programming language-style prompts.
@@ 5.5 Different Programming Languages
Interestingly, the $F_{1}$ scores with CodeNER remain consistent between the Python and C++ instructions on CoNLL03. We think this suggests that dataset characteristics may have a more significant impact than the choice of the programming language. The comparable results between Python and C++ suggest that the choice of the programming language has a minimal impact on the effectiveness of CodeNER when used on datasets where the model’s existing capabilities are strong. We additionally considered using the Java programming language for CodeNER. However, the performance was nearly zero in terms of the F1 scores, which indicates LLMs lack inherent knowledge of Java in their pre-training Haroon et al. (2025).
@@ Appendix G Additional Case study
Dataset Template Type Phi-3 Llama-3 -8B Llama-3 -70B
FIN GoLLIE English 8.26 9.21 11.02
Translation 8.26 9.21 11.02
CodeNER English 17.51 31.96 19.39
Translation 17.51 31.96 19.39
w/o label 8.49 45.62 26.11
CoNLL03 GoLLIE English 35.39 34.72 53.93
Translation 35.39 34.72 53.93
CodeNER English 36.57 40.29 54.94
Translation 36.57 40.29 54.94
w/o label 48.33 37.00 47.53
SwissNER GoLLIE English 32.54 24.98 44.49
Translation 27.59 10.04 39.45
CodeNER English 32.16 35.03 49.12
Translation 28.76 35.48 47.04
w/o label 31.78 31.87 39.21
Arabic A GoLLIE English 14.60 12.57 22.83
Translation 3.07 7.57 21.72
CodeNER English 9.18 16.24 29.47
Translation 5.04 20.42 28.83
w/o label 8.81 21.84 29.00
Arabic B GoLLIE English 15.83 16.53 25.12
Translation 2.35 8.70 25.42
CodeNER English 9.31 16.58 33.78
Translation 6.79 27.52 32.16
w/o label 9.09 24.69 33.62
Finnish A GoLLIE English 23.48 21.86 34.15
Translation 19.61 6.58 26.72
CodeNER English 32.37 33.29 53.16
Translation 27.34 29.46 48.89
w/o label 31.69 32.11 48.98
DaNE GoLLIE English 27.37 25.33 47.17
Translation 21.82 17.11 43.68
CodeNER English 38.25 36.53 42.50
Translation 33.79 37.55 42.93
w/o label 41.27 30.02 35.06
Average GoLLIE English 22.50 20.74 34.10
Translation 16.87 13.42 31.71
CodeNER English 25.05 29.99 40.34
Translation 22.26 31.81 39.17
w/o label 25.64 31.88 37.07
Table 11: Experimental macro F1-scores for GoLLIE and CodeNER using label descriptions either translated into each dataset’s language or kept in English. In the type column, Translation denotes labels translated with Google Translator, while English indicates the original prompt remained unchanged.